OpenStudy (anonymous):

Gary is using a combination of regular polygonal tiles for his kitchen floor. He begins the tessellation at a vertex by placing one hexagon and two triangles. Which two polygons should he use to fill up the vacant space at the vertex and make it possible to tessellate the floor? A)2 squares because 180 = 90 + 90 B)1 triangle and 1 hexagon because 180 = 60 + 120 c)2 hexagons because 360 – (120 + 60 +60) = 120 d)2 triangles because 360 – (120 + 60 +60) = 2 x 60

OpenStudy (hba):

t depends on the angle (or angles) the two triangles make at that vertex. What is certainly true is the the sum of all the angles of the figures at that vertex must equal 360 degrees. But we cannot know what other figures to place there without knowing more about the first two triangles.

OpenStudy (anonymous):

hmm well thats all they gave me

OpenStudy (hba):

u have 4 days for this assignment right

OpenStudy (anonymous):

i only have this 2 more problems then im done(:

OpenStudy (hba):

k

OpenStudy (anonymous):

1 square and a hexagon because 360 – (60 + 60 + 60) = 60 + 120 1 triangle and a hexagon because 180 = 60 + 120 2 hexagons because 360 – (60 + 60) = 2 x 120 2 squares because 180 = 90 + 90

OpenStudy (anonymous):

So a is ur answer

OpenStudy (anonymous):

Well, (d) for two polygons that fit in the 120 degree gap at the vertex.?

OpenStudy (anonymous):

and 1 square and a hexagon are not a choice...

OpenStudy (anonymous):

well yes but when you draw it out you can see how many squares fit in it and if you check the rest youll see it wont work.

OpenStudy (anonymous):

but what you said isnt a option : A)2 squares because 180 = 90 + 90 B)1 triangle and 1 hexagon because 180 = 60 + 120 c)2 hexagons because 360 – (120 + 60 +60) = 120 d)2 triangles because 360 – (120 + 60 +60) = 2 x 60

OpenStudy (anonymous):

A)2 squares because 180 = 90 + 90

OpenStudy (anonymous):

okay whatever you want too choose your the worker not us.

OpenStudy (anonymous):

okay whatever you want too choose your the worker not us.

OpenStudy (anonymous):

im aware/

OpenStudy (anonymous):

okay whatever you want too choose your the worker not us.

OpenStudy (anonymous):

why say it 3 times?

OpenStudy (anonymous):

ok sorry i got it all wrong. Well, answer is (d) for two polygons that fit in the 120 degree gap at the vertex.

OpenStudy (anonymous):

yay (: